# Exceptional Tannaka groups only arise from cubic threefolds

### v2

@@ 4.1. Reduction to semisimple groups
from section 3.1 satisfies Hodge symmetry in the sense that $\dim V^{p,q}=\dim V^{q,p}$ for all $p,q\in\mathbb{Z}$. Recall that this decomposition is given by the weights for the Hodge cocharacter
@@ 4.2. Cocharacters with Hodge properties
be the set of characters that enter in the given representation. Condition (H2) then implies
$|\langle\lambda,\chi\rangle|\;\leqslant\;\ell\;<\;\dim V\quad\text{for all}\quad\chi\;\in\;S.$
This condition leaves only finitely many choices of $\lambda\in X_{\ast}(T)=\operatorname{Hom}(\mathbb{G}_{m},T)$ in the cocharacter lattice, since in the dual character lattice the occurring weights $\chi\in S$ span a subgroup
$\langle S\rangle_{\mathbb{Z}}\;\subset\;X^{\ast}(T)$
of finite index: Indeed, if the subgroup spanned by those weights had smaller rank than $X^{\ast}(T)$, then we could find $\mu\in X_{\ast}(T)\smallsetminus\{0\}$ with $\langle\mu,\chi\rangle=0$ for all $\chi\in S$. But then the image of
$\mu\colon\quad\mathbb{G}_{m}\;\longrightarrow\;G$
would be a subgroup of positive dimension acting trivially on the representation $V$ which is impossible since $V$ has finite kernel. ∎
@@ 4.3. Cocharacters of $G=E_{6}$
Consider the simply connected group $G=E_{6}$. Fix a system of positive roots for a maximal torus $T\subset G$. We label the fundamental dominant weights
$\varpi_{1},\dots,\varpi_{6}\;\in\;X^{\ast}(T)$
as in [Bou02, ch. VI, sect. 4.12], and the dual basis by $\varpi_{1}^{\vee},\dots,\varpi_{6}^{\vee}\in X_{\ast}(T)$. An exhaustive computer search gives:
@@
Let $V$ be a $27$-dimensional irreducible representation of $E_{6}$, then for any cocharacter
$\lambda\colon\quad\mathbb{G}_{m}\;\longrightarrow\;E_{6}$
with (H1), (H2), (H3), the Hodge numbers $h^{i}(\lambda):=\dim V^{2i-\ell}(\lambda)$ for $0\leqslant i\leqslant\ell$ are given by one of the rows of the following table:
$\begin{array}[]{c|c|c|c|c|c|c}h^{0}&h^{1}&h^{2}&h^{3}&h^{4}&h^{5}&h^{6}\\ \hline\cr 6&15&6&&&&\\ 3&6&9&6&3&&\\ 3&3&3&9&3&3&3\\ \end{array}\vskip 6.0pt plus 2.0pt minus 2.0pt$
Moveover, up to conjugacy the only cocharacter giving the Hodge numbers $6,15,6$ in the first row of the table is the cocharacter $\lambda=\varpi_{1}^{\vee}$.
@@
$\dim W^{2i-\ell}(\lambda)\;=\;\begin{cases}1&\text{for\leavevmode\nobreak\ $i=0,4$},\\ 20&\text{for\leavevmode\nobreak\ $i=1,3$},\\ 36&\text{for\leavevmode\nobreak\ $i=2$}.\end{cases}$
This information will be useful in the proof of proposition 4.8 below.
@@ 4.4. Cocharacters of $G=E_{7}$
Let $V$ be a $56$-dimensional irreducible representation of $E_{7}$, then for any cocharacter
$\lambda\colon\quad\mathbb{G}_{m}\;\longrightarrow\;E_{7}$
with (H1), (H2), (H3) whose associated Hodge decomposition has odd length $\ell$, the corresponding Hodge numbers $h^{i}(\lambda)=\dim V^{2i-\ell}(\lambda)$ for $0\leqslant i\leqslant\ell$ are given by one of the rows of the following table:
$\begin{array}[]{c|c|c|c|c|c|c|c|c|c|c|c|c|c|c|c}h^{0}&h^{1}&h^{2}&h^{3}&h^{4}&h^{5}&h^{6}&h^{7}&h^{8}&h^{9}&h^{10}&h^{11}&h^{12}&h^{13}&h^{14}&h^{15}\\ \hline\cr 7&21&21&7&&&&&&&&&&&&\\ 6&7&15&15&7&6&&&&&&&&&&\\ 6&6&1&15&15&1&6&6&&&&&&&&\\ 5&3&10&10&10&10&3&5&&&&&&&&\\ 5&2&6&5&10&10&5&6&2&5&&&&&&\\ 4&3&3&12&6&6&12&3&3&4&&&&&&\\ 5&2&5&1&5&10&10&5&1&5&2&5&&&&\\ 4&2&3&5&8&6&6&8&5&3&2&4&&&\\ 4&2&2&5&1&8&6&6&8&1&5&2&2&4&\\ 4&2&2&4&1&1&8&6&6&8&1&1&4&2&2&4\\ \end{array}\vskip 6.0pt plus 2.0pt minus 2.0pt$
@@ 4.5. Hodge numbers of exceptional subvarieties
    $G_{X,\omega}^{\ast}\simeq E_{7}$$\dim V_{X}=56$ and $d\in\{3,\dots,15\}$ is odd.
@@
Since $X\subset A$ is a smooth subvariety, the Tannaka group $G_{X,\omega}^{\ast}$ acts on $V_{X}$ via a minuscule representation [Krä22, cor. 1.10]. For the simple exceptional groups the only minuscule representations are the irreducible representations of dimension $27$ for $E_{6}$ and of dimension $56$ for $E_{7}$. The latter representation admits a nondegenerate bilinear form which is alternating, so in this case $X$ must have odd dimension by [JKLM23, sect. 1.2]; hence for $E_{7}$ we only need to consider cocharacters whose associated Hodge decomposition has odd length $d$. Also, when $X$ is a divisor, the topological Euler characteristic of $X$ is divisible by $g!$, hence cannot be $27$ or $56$ up to sign. Therefore, we can assume $d\leqslant g-2$. Since $X\subset A$ has ample normal bundle, lemma 3.4 and corollary 3.5 imply that the Hodge cocharacter induces a cocharacter of $G_{X,\omega}^{\ast}$ satisfying properties (H1), (H2) and (H3). Lemma 3.4 also implies the following bound on the codimension:
@@
So it follows that the Hodge numbers $h^{p}(X)=(-1)^{d-p}\chi(X,\Omega^{p}_{X})$ must be among those in propositions 4.3, LABEL: and 4.5. ∎
@@ 4.6. Surfaces with group $E_{6}$ and the difference morphism
Let us take a closer look at the first row of the table in corollary 4.7. The corollary in particular says that if $X\subset A$ is a smooth nondivisible irreducible subvariety with ample normal bundle and dimension $d<g/2$ such that $G_{X,\omega}^{\ast}\simeq E_{6}$, then $X$ must be a surface with
@@
We also know the Chern numbers: The top Chern class is $c_{2}(X)=\dim\omega(\delta_{S})=27$, hence $c_{1}(X)^{2}=45$ by Noether’s formula. To show that $X$ is indeed the Fano surface of lines on a smooth cubic threefold, we will need one more piece of numerical information: The degree of the difference morphism
@@
Let $D:=X-X$. Since $X$ is nondegenerate, the morphism $d\colon X\times X\to D$ is generically finite, hence over some open dense subset $U\subset D^{\textup{reg}}$ it restricts to a finite étale cover of degree $\deg(d)$. The direct image $\delta_{X}*\delta_{-X}=Rd_{*}(\delta_{X\times X})$ therefore restricts on $U$ to a local system of rank $\deg(d)$, placed in cohomological degree $-\dim D=-4$. By adjunction that local system contains the trivial local system of rank one as a direct summand. The decomposition theorem therefore shows that
@@
where $P$ is a semisimple perverse sheaf on $D$ which restricts on $U\subset D$ to a local system of rank $\deg(d)-1$, placed in cohomological degree $-4$. Passing to the clean characteristic cycles of these perverse sheaves in the sense of [JKLM23, sect. 5.3], we get
@@
To interpret this in terms of representation theory, we need some notation. Fix a maximal torus $T\subset E_{6}$. We use the fundamental weights $\varpi_{1},\dots,\varpi_{6}\in X^{\ast}(T)$ from section 4.3. The dominant integral weights are precisely those of the form
@@
In these terms the representation ring $\textup{R}(E_{6})$ has a basis consisting of the Weyl group orbits of the dominant integral weights. For a weight $\lambda\in X^{\ast}(T)$ we denote by $[\lambda]\in\textup{R}(E_{6})$ its Weyl group orbit. The character of a representation is a single Weyl group orbit only if the representation is minuscule, so the basis given by Weyl group orbits differs from the one given by the irreducible representations. In what follows we are interested in the irreducible representations that appear in the tensor product of the 27-dimensional representation $V=V_{\varpi_{1}}$ with its dual $V^{\vee}=V_{\varpi_{6}}$:
(2a) $V\otimes V^{\vee}\;\simeq\;V_{\alpha}\oplus V_{\beta}\oplus\bf{1}\quad\text{with}\quad\begin{cases}\alpha\;=\;\varpi_{2},\\ \beta\;=\;\varpi_{1}+\varpi_{6}.\end{cases}$
@@
Now we come back to geometry. Fix an isomorphism $G_{X,\omega}^{\ast}\simeq E_{6}$. By corollary 2.5 we have $G_{\omega}^{\ast}(\delta_{X}^{H})=G_{X,\omega}^{\ast}$, so we may work as well with Hodge modules. Since $\omega(\delta_{X}^{H})$ is a minuscule representation, it must be one of the two irreducible representations of dimension 27, which are dual to each other. So we may assume $\omega(\delta_{X}^{H})\simeq V$. Then eq. 2a translates to
@@
where for $\lambda\in\{\alpha,\beta\}$, we denote by $M_{\lambda}\in\langle\delta_{X}^{H}\rangle$ the simple Hodge module with associated representation $\omega(M_{\lambda})\simeq V_{\lambda}$. The characteristic cycles of these Hodge modules can be computed as follows: By [JKLM23, th. 5.11, lemma 5.14] there is a natural way to attach to any Weyl group orbit $[\lambda]$ an effective Lagrangian cycle $\Lambda_{\lambda}$ on the cotangent bundle $T^{*}A$ such that if we extend this map additively to the representation ring $\textup{R}(E_{6})$, then it sends the character of any representation to the characteristic cycle of the corresponding perverse sheaf. Thus eqs. 2b, LABEL: and 2c translate to
@@
Comparing eq. 1a and eq. 3a, we see that $\delta_{D}^{H}\simeq M_{\lambda}$ for some $\lambda\in\{\alpha,\beta\}$. We claim that
@@
Indeed, if $\delta_{D}^{H}\simeq M_{\beta}$, then $\operatorname{cc}(M_{\beta})$ would contain the irreducible component $\Lambda_{D}$ with multiplicity one. By eq. 3c this component would then have to enter in the summand $\Lambda_{\beta}$, since the other summands enter with higher multiplicity. Then by eq. 1b the other summand $\Lambda_{\alpha}$ could involve only conormal varieties $\Lambda_{Z}$ with
@@
But then by eq. 3b the cycle $\operatorname{cc}(M_{\alpha})$ would also only involve conormal varieties $\Lambda_{Z}$ with $\dim(Z)<4$. This would imply
[…]
hence $\dim d^{-1}(\operatorname{Supp}M_{\alpha})<4$ because $d\colon X\times X\to D$ is dominant and $X\times X$ is irreducible. By lemma 3.7 then
[…]
But $\omega(M_{\alpha})\simeq V_{\alpha}$ is the adjoint representation of $E_{6}$, and by remark 4.4 the Hodge cocharacter induces on the adjoint representation a Hodge decomposition of level $4$, a contradiction. This proves that $\delta_{D}^{H}\simeq M_{\alpha}$ as claimed. But then by eq. 3b the conormal variety $\Lambda_{D}$ enters as a component in the cycle $\Lambda_{\alpha}$. It then follows that the cycle
[…]
@@
The above numerical results in fact characterize Fano surfaces of smooth cubic threefolds, as we will see in the classification of subvarieties with Tannaka group $E_{6}$ in theorem 7.1. In particular, the degree of the morphism $d\colon X\times X\to X-X$ is precisely $6$. However, for this will not use any representation theory; instead, we will show by direct geometric arguments that any smooth projective surface with the above numerical properties is isomorphic to the Fano surface of a smooth cubic threefold. As a preparation for this, we gather in the next two sections some general facts about cubic hypersurfaces and Fano varieties of lines.

### v3

@@ 4.1. Reduction to semisimple groups
from section˜3.1 satisfies Hodge symmetry in the sense that $\dim V^{p,q}=\dim V^{q,p}$ for all $p,q\in\mathbb{Z}$. Recall that this decomposition is given by the weights for the Hodge cocharacter
@@ 4.2. Cocharacters with Hodge properties
be the set of characters that enter in the given representation. Condition (H2) then implies $|\langle\lambda,\chi\rangle|\leqslant\ell<\dim V$ for all $\chi\in S$. This condition leaves only finitely many choices of $\lambda\in X_{\ast}(T)=\operatorname{Hom}(\mathbb{G}_{m},T)$ in the cocharacter lattice, since in the dual character lattice the occurring weights $\chi\in S$ span a subgroup $\langle S\rangle_{\mathbb{Z}}\subset X^{\ast}(T)$ of finite index: Indeed, if the subgroup spanned by those weights had smaller rank than $X^{\ast}(T)$, then we could find $\mu\in X_{\ast}(T)\smallsetminus\{0\}$ with $\langle\mu,\chi\rangle=0$ for all $\chi\in S$. But then the image of $\mu\colon\mathbb{G}_{m}\to G$ would be a subgroup of positive dimension acting trivially on the representation $V$ which is impossible since $V$ has finite kernel. ∎
@@ 4.3. Cocharacters of $G=E_{6}$
Consider the simply connected group $G=E_{6}$. Fix a system of positive roots for a maximal torus $T\subset G$. We label the fundamental dominant weights $\varpi_{1},\dots,\varpi_{6}\in X^{\ast}(T)$ as in [Bou02, ch. VI, sect. 4.12], and the dual basis by $\varpi_{1}^{\vee},\dots,\varpi_{6}^{\vee}\in X_{\ast}(T)$. An exhaustive computer search gives:
@@
Let $V$ be a $27$-dimensional irreducible representation of $E_{6}$, then for any $\lambda\in X_{\ast}(E_{6})$ with (H1), (H2), (H3), the numbers $h^{i}(\lambda):=\dim V^{2i-\ell}(\lambda)$ for $0\leqslant i\leqslant\ell$ are given by one of the rows of the following table:
[…]
Moreover, up to conjugacy the only cocharacter giving the Hodge numbers $6,15,6$ in the first row of the table is the cocharacter $\lambda=\varpi_{1}^{\vee}$.
@@
$\dim W^{2i-\ell}(\lambda)\;=\;\begin{cases}1&\text{for\penalty 10000\ $i=0,4$},\\ 20&\text{for\penalty 10000\ $i=1,3$},\\ 36&\text{for\penalty 10000\ $i=2$}.\end{cases}$
This information will be useful in the proof of proposition˜4.8 below.
@@ 4.4. Cocharacters of $G=E_{7}$
Let $V$ be a $56$-dimensional irreducible representation of $E_{7}$, then for any $\lambda\in X_{\ast}(E_{7})$ with (H1), (H2), (H3) whose associated Hodge decomposition has odd length $\ell$, the numbers $h^{i}(\lambda)=\dim V^{2i-\ell}(\lambda)$ for $0\leqslant i\leqslant\ell$ are given by one of the rows of the following table:
[…]
@@ 4.5. Hodge numbers of exceptional subvarieties
    $G_{X,\omega}^{\ast}\simeq E_{7}$, $\dim V_{X}=56$ and $d\in\{3,\dots,15\}$ is odd.
@@
Since $X\subset A$ is a smooth subvariety, the Tannaka group $G_{X,\omega}^{\ast}$ acts on $V_{X}$ via a minuscule representation [Krä22, cor. 1.10]. For the simple exceptional groups the only minuscule representations are the irreducible representations of dimension $27$ for $E_{6}$ and of dimension $56$ for $E_{7}$. The latter representation admits a nondegenerate bilinear form which is alternating, so in this case $X$ must have odd dimension by [JKLM25, sect. 1.2]; hence for $E_{7}$ we only need to consider cocharacters whose associated Hodge decomposition has odd length $d$. Also, when $X$ is a divisor, the topological Euler characteristic of $X$ is divisible by $g!$, hence cannot be $27$ or $56$ up to sign. Therefore, we can assume $d\leqslant g-2$. Since $X\subset A$ has ample normal bundle, lemma˜3.4 and corollary˜3.5 imply that the Hodge cocharacter induces a cocharacter of $G_{X,\omega}^{\ast}$ satisfying properties (H1), (H2) and (H3). Lemma˜3.4 also implies the following bound on the codimension:
@@
So it follows that the Hodge numbers $h^{p}(X)=(-1)^{d-p}\chi(X,\Omega^{p}_{X})$ must be among those in propositions˜4.3, LABEL: and 4.5. ∎
@@ 4.6. Surfaces with group $E_{6}$ and the difference morphism
Let us take a closer look at the first row of the table in corollary˜4.7. The corollary in particular says that if $X\subset A$ is a smooth nondivisible irreducible subvariety with ample normal bundle and dimension <g/2<g/2 such that $G_{X,\omega}^{\ast}\simeq E_{6}$, then $X$ must be a surface with
@@
We also know the Chern numbers: The top Chern class is $c_{2}(X)=\dim\omega(\delta_{S})=27$, so $c_{1}(X)^{2}=45$ by Noether’s formula. Moreover, by the representation theory of $E_{6}$ there is a unique one-dimensional subrepresentation inside $V_{X}\otimes V_{X}\otimes V_{X}$. Since by [Wei15, sect. 10] or [Krä22, lemma 3.5] the one-dimensional representations of the Tannaka group correspond to skyscraper sheaves, it follows that there a unique skyscraper direct summand inside $\delta_{X}*\delta_{X}*\delta_{X}$. By base change this means that the sum morphism $X^{3}\to A$ has a unique three-dimensional fiber and that this fiber is irreducible. To show that $X$ is the Fano surface of lines on a smooth cubic threefold (in which case the above fiber parametrizes coplanar triples of lines), we will need one more piece of numerical information: The degree of the difference morphism
@@
Let D:=XXD:=X-X. Since $X$ is nondegenerate, the morphism $d\colon X\times X\to D$ is generically finite, hence over some open dense subset $U\subset D^{\textup{reg}}$ it restricts to a finite étale cover of degree $\deg(d)$. The direct image $\delta_{X}*\delta_{-X}=Rd_{*}(\delta_{X\times X})$ therefore restricts on $U$ to a local system of rank $\deg(d)$, placed in cohomological degree $\dim D=-4$. By adjunction that local system contains the trivial local system of rank one as a direct summand. The decomposition theorem [BBDG18, th. 6.2.5] therefore shows that
@@
where $P$ is a semisimple perverse sheaf on $D$ which restricts on $U\subset D$ to a local system of rank $\deg(d)-1$, placed in cohomological degree 4-4. Passing to the clean characteristic cycles of these perverse sheaves in the sense of [JKLM25, sect. 5.3], we get
@@
To interpret this in terms of representation theory, we need some notation. Fix a maximal torus $T\subset E_{6}$. We use the fundamental weights $\varpi_{1},\dots,\varpi_{6}\in X^{\ast}(T)$ from section˜4.3. The dominant integral weights are precisely those of the form
@@
In these terms the representation ring $\textup{R}(E_{6})$ has a basis consisting of the Weyl group orbits of the dominant integral weights. For a weight $\lambda\in X^{\ast}(T)$ we denote by $\lambda]\in\textup{R}(E_{6})$ its Weyl group orbit. The character of a representation is a single Weyl group orbit only if the representation is minuscule, so the basis given by Weyl group orbits differs from the one given by the irreducible representations. In what follows we are interested in the irreducible representations that appear in the tensor product of the 27-dimensional representation $V=V_{\varpi_{1}}$ with its dual $V^{\vee}=V_{\varpi_{6}}$:
(2a) $V\otimes V^{\vee}\;\simeq\;V_{\alpha}\oplus V_{\beta}\oplus\bf{1}\qquad\text{with}\qquad\begin{cases}\alpha\;=\;\varpi_{2},\\ \beta\;=\;\varpi_{1}+\varpi_{6}.\end{cases}$
@@
Now we come back to geometry. Fix an isomorphism $G_{X,\omega}^{\ast}\simeq E_{6}$. By corollary˜2.5 we have $G_{\omega}^{\ast}(\delta_{X}^{H})=G_{X,\omega}^{\ast}$, so we may work as well with Hodge modules. Since $\omega(\delta_{X}^{H})$ is a minuscule representation, it must be one of the two irreducible representations of dimension 27, which are dual to each other. So we may assume $\omega(\delta_{X}^{H})\simeq V$. Then eq.˜2a translates to
@@
where for $\lambda\in\{\alpha,\beta\}$, we denote by $M_{\lambda}\in\langle\delta_{X}^{H}\rangle$ the simple Hodge module with associated representation $\omega(M_{\lambda})\simeq V_{\lambda}$. The characteristic cycles of these Hodge modules can be computed as follows: By [JKLM25, th. 5.11, lemma 5.14] there is a natural way to attach to any Weyl group orbit $\lambda]$ an effective Lagrangian cycle $\Lambda_{\lambda}$ on the cotangent bundle $T^{*}A$ such that if we extend this map additively to the representation ring $\textup{R}(E_{6})$, then it sends the character of any representation to the characteristic cycle of the corresponding perverse sheaf. Thus eqs.˜2b, LABEL: and 2c translate to
@@
Comparing eq.˜1a and eq.˜3a, we see that $\delta_{D}^{H}\simeq M_{\lambda}$ for some $\lambda\in\{\alpha,\beta\}$. We claim that
@@
Indeed, if $\delta_{D}^{H}\simeq M_{\beta}$, then $\operatorname{cc}(M_{\beta})$ would contain the irreducible component $\Lambda_{D}$ with multiplicity one. By eq.˜3c this component would then have to enter in the summand $\Lambda_{\beta}$, since the other summands enter with higher multiplicity. Then by eq.˜1b the other summand $\Lambda_{\alpha}$ could involve only conormal varieties $\Lambda_{Z}$ with
@@
But then by eq.˜3b the cycle $\operatorname{cc}(M_{\alpha})$ would also only involve conormal varieties $\Lambda_{Z}$ with $\dim(Z)<4$. This would imply
$\dim\operatorname{Supp}M_{\alpha}\;<\;\dim\operatorname{Supp}D\;=\;4,$
hence $\dim d^{-1}(\operatorname{Supp}M_{\alpha})<4$ because $d\colon X\times X\to D$ is dominant and $X\times X$ is irreducible. By lemma˜3.7 then
$h^{p,q}(M_{\alpha})\;=\;0\quad\text{for}\quad|p-q|\;>\;3.$
But $\omega(M_{\alpha})\simeq V_{\alpha}$ is the adjoint representation of $E_{6}$, and by remark˜4.4 the Hodge cocharacter induces on the adjoint representation a Hodge decomposition of level $4$, a contradiction. This proves that $\delta_{D}^{H}\simeq M_{\alpha}$ as claimed. But then by eq.˜3b the conormal variety $\Lambda_{D}$ enters as a component in the cycle $\Lambda_{\alpha}$. It then follows that the cycle
$\operatorname{cc}(\delta_{X}^{H}*\delta_{-X}^{H})\;=\;26\cdot\Lambda_{0}+6\cdot\Lambda_{\alpha}+1\cdot\Lambda_{\beta}$
@@
The above numerical results in fact characterize Fano surfaces of smooth cubic threefolds, as we will see in the classification of subvarieties with Tannaka group $E_{6}$ in theorem˜7.1. In particular, the degree of the morphism $d\colon X\times X\to X-X$ is precisely $6$. However, for this will not use any representation theory; instead, we will show by direct geometric arguments that any smooth projective surface with the above numerical properties is isomorphic to the Fano surface of a smooth cubic threefold. As a preparation for this, we gather in the next two sections some general facts about cubic hypersurfaces and Fano varieties of lines.